OpenStudy (anonymous):

Please check my answer! What is the slope of the line that passes through the points (3, –4) and (–7, 1)? A.2 B.1/2 C.-2 <----- D. -1/2 <----- Either C or D

OpenStudy (anonymous):

Well, you're close, but you're incorrect. Do you know how to find slope? If not, I can help you through.

OpenStudy (anonymous):

i did that and got -5/10

OpenStudy (anonymous):

Think of rise and run.

OpenStudy (anonymous):

its 2 not -2...am i right?

OpenStudy (anonymous):

Hopefully this'll help.

OpenStudy (anonymous):

Your "rise" is your numerator, and your "run" is your denominator.

OpenStudy (anonymous):

Thanks can you help me with another?

OpenStudy (anonymous):

Sure!

OpenStudy (anonymous):

okay, hold on

OpenStudy (anonymous):

Which is the equation of the line? A. –5x – 3y = –15 B. 5x + 3y = 15 C. 5x – 3y = 15 D. –5x + 3y = 15

OpenStudy (anonymous):